Simon Companys year-end balance sheets follow. Current Yr 1 Yr Ago 2 Yrs Ago At December 31 Assets Cash Accounts receivable,Compute days sales uncollected. Days Sales Uncollected Choose Numerator: 1 Choose Denominator: Days = Days Sales UncollectFor each ratio, determine if it improved or worsened in the current year. Days sales uncollectedCompute accounts receivable turnover. Accounts Receivable Turnover 1 Choose Denominator: Choose Numerator: = Accounts ReceivaFor each ratio, determine if it improved or worsened in the current year........ Accounts receivable turnover
Answer #1

>> Average Accounts receivable for current year = ( $ 89,500 + $ 62,500 ) / 2 = $ 76,000.

>> Average Accounts receivable for 1 year Ago = ( $ 50,200 + $ 62,500 ) / 2 = $ 56,350.

___________________________________________________________________________

>> Days in sales Uncollected = ( Average Accounts receivable / Sales ) * 365 .

>> Days in sales Uncollected for current year = ( $ 76,000 / 673,500 ) * 365 = 41.19 Days.

>> Days in sales Uncollected for 1 year Ago = ( $ 56,350 / 532,000 ) * 365 = 38.66 Days.

_______________________________________________________________________________

>> Accounts receivable turnover = Sales / Average Accounts Receivable.

>> Accounts receivable turnover for current year = $ 673,500 / $ 76,000 = 8.862 times

>> Accounts receivable turnover for 1 year Ago = $ 532,000 / $ 56,350 = 9.44 times.

______________________________________________________________________________